OpenStudy (anonymous):

what is the vertex of the quadratic equation y=-2x^2 + 4x – 6??

OpenStudy (anonymous):

Use -b/2(a)

OpenStudy (anonymous):

so, -4/-2(2)

OpenStudy (anonymous):

Then once you get the value once you solves plug that number in for x and solve for y.

OpenStudy (anonymous):

From @orphtroll, Vertex ( h, k ): h = -b/2a = .... -> k = f (1) = -2 + 4 – 6 = ...